• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Boeglin

    Log entry time 16:12:12 on March13,2005

    Entry number 141387

    keyword=Shift Summary

    8:00 last run on LD : 2623 running
    8:15 finished LD change to Carbon (santed target)
    8:40 problems getting 40 uA. MCC tries to deliver 30 and investigate the high trip rate

    8:45 20 uA, start running with lower current
    9:00 get 40uA of beam. Start a new run with higher current

    8:45 start run 2624 (55K, current between 20 and 40 uA)
    9:04 start run 2625 (307K, current 40 uA, possible trip of NDET E, dE towards the end of the run)

    10:07 problem with HV for AUX, dE and E detectors. Crate 11 had to be reset manually. Lost communication with the crates. Possibly the comunnications VME does not boot right. Checked the HV's using the terminal servers. This required several accesses.

    12:50 tried to get beam. Had to reset Crate 12 (another access)

    13:03 decided to run since HV seems to be OK.

    run 2526 daq crashed. probably useless (very short)

    13:00 start 2627 (400 K, end failed)
    2628 (junk)
    do runs with 200K events to prevent scaler overflows

    14:28 start run 2629 (200 K)
    15:08 start run 2630 (200 K)
    15:50 start run 2631

    iochacs still not bootable. Check HV using the RS232 and watch the scalers